OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 3x + y = 14 –3x + 2y = 1 my answer is 5 Am I right?

OpenStudy (uri):

3x + y = 14 –3x + 2y = 1 ======== 3y=15 y=15/3 y=5 Correct!! :)

OpenStudy (anonymous):

Thank you! :)